# Description length of canonical and microcanonical models

### v2

@@ Abstract
Non-equivalence between the canonical and the microcanonical ensemble has been shown to arise for models defined by an extensive (i.e. ‘scaling with the size of the system’) number of constraints (e.g. the Configuration Model). Here, we compare the description length of binary canonical and microcanonical models in light of ensemble non-equivalence. Specifically, we consider the description length induced by the Normalized Maximum Likelihood (NML), which consists of two terms, i.e. a model log-likelihood and its complexity. While the effects of ensemble non-equivalence on the log-likelihood term are well understood, its effects on the complexity term have not yet been systematically studied. Here, we find that i) microcanonical models are always more complex than their canonical counterparts, ii) the best-scoring model in terms of description length highly depends on the numerical values of the constraints and iii) the difference between the canonical and the microcanonical description length is strongly influenced by the degree of non-equivalence, a result suggesting that non-equivalence should be taken into account when selecting models. Finally, we compare the NML-based approach to model selection with the Bayesian one in light of our results, showing that the Bayesian description length becomes much more sensitive to the choice of the prior when an extensive number of constraints is involved.
@@ I Introduction
Entropy maximization provides a principled approach to inference, resulting in models that are maximally random given certain constraints. These models have found applications across a wide range of scientific disciplines, such as network science squartini_springer ; cimini19 ; squartini18 and time-series analysis marcaccioli2020maximum ; marcaccioli2020correspondence .
@@
Two distinct formulations of maximum-entropy models are possible depending on how constraints are enforced. Canonical models are defined by requiring the value of the constraints to be satisfied on average. Microcanonical models are defined by requiring the value of the constraints to be satisfied exactly. These two formulations do not necessarily lead to the same description of a given system, i.e. they are, in general, non-equivalent.
@@
Non-equivalence between canonical and microcanonical models can be formally defined in several (under mild conditions, equivalent) ways touchette15 . A particularly convenient definition is the measure-level one, which is rooted in information theory and involves the Kullback-Leibler (KL) divergence between the microcanonical and canonical probability distributions. According to this definition, (non-)equivalence is characterized by a (non-)vanishing relative entropy density as the system size approaches infinity. Interestingly, since the KL divergence coincides with the difference between the canonical and the microcanonical log-likelihood functions, measure-level (non-)equivalence can be inspected by simply considering the asymptotic behaviour of such difference squartini15 . Non-equivalence has been recently shown to arise in models that involve an extensive number of constraints (e.g. the Configuration Model for networks squartini15 ; squartini17 ; qi ) and may have practical implications for the expected macroscopic properties of the system under study (see bruno20 for an example concerning bipartite networks).
@@
Whenever the two approaches are equivalent, they identify the same set of typical configurations. In these cases, preferring one formulation over the other becomes a matter of convenience (e.g. computational efficiency). Whenever the two approaches are non-equivalent, the choice depends on the specific characteristics of the system under study and the nature of the constraints involved. For example, in the context of choosing a null model to be compared with real data for e.g. pattern detection, it has been argued that, when non-equivalence holds, the ensemble choice should be based on a theoretical guiding principle garlaschelli17 ; qi ; squartini15bis ; garlaschelli18 ; garlaschelli16 . In this case, if the data are expected to be noisy, meaning that the measured values of the constraints generally do not match the true (unobservable) ones, one should prefer the canonical version of the null model because it allows for fluctuations in the constraint quantities, while the microcanonical one would paradoxically assign zero probability to the true configuration of the system. Conversely, if, by hypothesis, the measured values of the constraints are not affected by error or noise, the microcanonical model should be preferred.
@@
When there is no prior expectation available about the presence or absence of noise in the empirical values of the constraints, the decision between hard and soft constraints should be based on posterior evidence, i.e. which of the two null models best describes the data. This immediately leads to a problem of model selection between canonical and microcanonical ensembles.
@@
Information theory provides a powerful framework for model selection by offering criteria that capture the trade-off between a model ‘goodness of fit,’ quantified by its log-likelihood, and a model ‘complexity.’ The two historically most popular model selection criteria are the Akaike information criterion (AIC) aic and the Bayesian information criterion (BIC) bic , according to which one should pick the model minimizing
@@
respectively, where $\mathcal{L}$ is the model likelihood, $k$ is the number of its parameters and $V$ is the sample size. Canonical and microcanonical models defined by the same set of constraints share the same number of parameters. Therefore, according to both AIC and BIC, their complexity is the same, and a comparison based on these criteria would reduce to a mere comparison between their log-likelihood functions. This would naively lead to the conclusion that the microcanonical model is always the best-scoring one, since it has the highest likelihood. However, neither AIC nor BIC can be computed for microcanonical models, since they are derived under regularity assumptions that are not even remotely met by the latter. Moreover, both criteria are asymptotic results derived under the assumption of a finite number of parameters as the system size approaches infinity burnham . This assumption is not met by models showing non-equivalence, whose number of parameters scales with the size of the system.
@@
To overcome these issues, here we consider the Minimum Description Length (MDL) principle grunwald_book ; grunwald2019 ; yamanishi2023 , a family of information theoretical approaches to model selection seeking the model that provides the shortest description of data. Specifically, we focus on the approach based upon the Normalized Maximum Likelihood (NML), or Shtarkov, distribution rissanen96 , our aim being that of investigating the interplay between the difference of log-likelihood functions and that of complexities in light of ensemble non-equivalence. Indeed, while the impact of non-equivalence on the difference between log-likelihoods is quite well understood, its effects on the difference between complexities remain an unexplored topic within this framework. To encompass a broad spectrum of applications, our study focuses on maximum-entropy models suitable for studying systems represented by binary rectangular matrices (e.g., bipartite networks, time series). More precisely, we aim at determining i) whether the influence of non-equivalence persists even when the complexity term is accounted for and ii) if non-equivalent canonical and microcanonical models should, in fact, be considered as different from a model selection perspective. We conclude that the answer to both questions is yes, raising the question of whether non-equivalence can be defined on a model selection level. Moreover, as a consequence of our results, we show that, in a Bayesian context, different priors can result in vastly different description lengths when an extensive number of constraints is enforced.
@@
The rest of the paper is organized as follows. Section II introduces MDL for discrete data and the Normalized Maximum Likelihood (NML) approach to define description lengths. In Section III, we provide a general expression for the difference between the description length of canonical and microcanonical models as a function of the Kullback-Leibler divergence between the two distributions. In Section IV, we explicitly derive and compare the description length of both the canonical and the microcanonical formulations of models defined by one global constraint (i.e. the sum of the entries of a matrix) as well as one-sided local constraints (i.e. the row-specific sums of a matrix). While the first constraint leads to ensemble equivalence, the second set of constraints leads to ensemble non-equivalence. Finally, Section V reviews the relationship between the NML-based and Bayesian approaches when non-equivalence holds.
@@ II The Minimum Description Length principle
However, the distribution $\mathcal{L}_{\mathcal{M}}(\bm{x})$ is affected by (at least) two problems having the same origin: the parameters appearing in its definition are evaluated for each individual data in $\mathcal{X}$. As a consequence, i) $\mathcal{L}_{\mathcal{M}}(\bm{x})$ cannot be a probability distribution with support in $\mathcal{X}$ as it is not properly normalized, and ii) adopting $\mathcal{L}_{\mathcal{M}}(\bm{x})$ may lead to severe overfitting as it is defined a posteriori with respect to the observations.
To solve these problems, one can build a so-called universal distribution relative to $\mathcal{M}$, i.e. a unique, representative distribution $\bar{P}_{\mathcal{M}}(\bm{x})$ of the statistical model $\mathcal{M}$. Once equipped with a universal distribution, the description length of $\bm{x}$ through is $\mathcal{M}$ is defined as
@@
Universal distributions (which don’t necessarily belong to $\mathcal{M}$) can be defined in several ways. In what follows, we will consider the MDL recipe based upon the NML universal distribution (for a more formal introduction to universal coding, we redirect the interested reader to grunwald_book ).
@@ II.1 The Normalized Maximum Likelihood universal distribution
The NML distribution corresponds to the distribution attaining the minimum point-wise regret in the worst-case scenario grunwald_book , i.e.
@@
In words, it is the distribution that a priori requires the minimum number of nats when compared to the shortest description length a posteriori, that is $-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))$. The minimum is found in the worst-case scenario, i.e. for the data $\bm{x}$ for which this difference is maximal. The unique solution to this minimax problem reads
@@
is named parametric complexity of $\mathcal{M}$. Notice that $\mathcal{X}$ represents the set of all possible data sets of fixed size $V$, i.e. all the observable samples of size $V$. Thus, $\text{COMP}_{\mathcal{M}}$ depends on both the model and $\mathcal{X}$. Finally, we can rewrite equation (II.1) as
@@ III Description length of canonical and microcanonical models
Although computing the complexity term within the NML-based description length can be challenging, models defined by a sufficient statistic admit a simpler expression for this term. This section introduces an important class of models of this kind, the maximum-entropy models (MEMs), both in their canonical and microcanonical formulations. Specifically, we will focus on discrete data that can be represented as a binary $n\times m$ matrix $\mathbf{G}$, i.e.
@@
and denote by $\mathcal{G}$ the set of all such matrices that, when endowed with a probability distribution $P(\mathbf{G})$, constitutes a proper ensemble of matrices. The matrix representation is often used to describe systems composed of $n$ elements to be modelled (corresponding to the number of rows), each characterized by $m$ state variables or degrees of freedom (corresponding to the number of columns). The size of the data set is fixed and supposed to represent the number of independent entries of the matrix, i.e. $nm$.
@@
while constraining the sufficient statistics. If the model is required to reproduce the value of the constraints exactly, i.e. $\mathbf{c}(\mathbf{G})=\mathbf{c}^{*}$, one obtains a microcanonical model. If, instead, the model is required to reproduce the value of the constraints on average, i.e. $\langle\mathbf{c}(\mathbf{G})\rangle_{P}=\mathbf{c}^{*}$ with $\langle\cdot\rangle_{P}$ being the ensemble average induced by distribution $P$, one obtains the corresponding canonical formulation. This approach, introduced by Gibbs gibbs and reprised by Jaynes jaynes , leads to ensembles of matrices that reproduce (either exactly or on average) the constrained quantities while randomizing everything else.
@@
with $\delta_{i,j}$ being the Kronecker delta. As this set may be empty for some values of $\mathbf{c}$, we will solely focus on the graphical values, i.e. the set $\mathcal{C}=\{\mathbf{c}:\Omega(\mathbf{c})\neq\emptyset\}$ of values that are verified by at least one configuration, and denote it with $\mathcal{N}_{\mathcal{C}}=|\mathcal{C}|$ its cardinality. In other words, the sum in equation (14) runs over the graphical values of the sufficient statistics rather than over all possible data.
@@ III.1 Description length of microcanonical models
$P_{\textrm{mic}}(\mathbf{G};\mathbf{c})=\begin{dcases}\frac{1}{\Omega(\mathbf{c})}&\text{if}\>\mathbf{c}(\mathbf{G})=\mathbf{c}\\ 0&\text{else},\end{dcases}$ (16)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{c})=\ln\frac{1}{\Omega(\mathbf{c})}=-\ln\Omega(\mathbf{c})$ (17)
@@ III.2 Description length of canonical models
In fact, it can be shown squartini17 ; squartini15 that
@@ III.3 Comparing canonical and microcanonical models
i.e. the canonical likelihood is always smaller than or equal to the microcanonical one. For the same reason, it is straightforward to see that the canonical complexity is always smaller than, or equal to, the microcanonical one:
@@
where $Q_{\text{can}}^{*}\equiv Q_{\text{can}}(\mathbf{c}^{*},\bm{\theta}^{*})$. This expression depends only on the canonical model, providing a new interpretation of the comparison between the canonical and the microcanonical description length: the canonical model provides a shorter description length whenever $\Delta\text{DL}(\mathbf{G}^{*})<0$, i.e. whenever $Q_{\text{can}}^{*}>\bar{Q}_{\text{can}}$. In other words, the canonical model is to be preferred over the microcanonical one whenever the maximum likelihood it assigns to the observed value $\mathbf{c}^{*}$ is higher than the maximum likelihood averaged over all possible values of the sufficient statistic (or, more intuitively, if it performs ‘better than average’ on the observed sufficient statistic). Remarkably, this is in open contrast with a comparison based only on the likelihoods, according to which the microcanonical formulation should always be preferred, independently of the observed data.
Expression (III.3) shows that the entity of the difference between the two description lengths crucially depends on the value of the KL divergence. As said in the introduction, the canonical and the microcanonical ensembles are (non-)equivalent if the relative $\text{D}_{\text{KL}}$ is (non-)vanishing as the system size approaches infinity. When studying matrices, identifying the system size with the number of entries, i.e. $nm$, may be inappropriate as adding (only) one entry to a matrix without altering its structure is, in fact, impossible. In these cases, it may be much more reasonable to identify the system size with the number $n$ of items: increasing the system size from $n$ to $n+1$ would, thus, correspond to adding a node to a network or a time-point to a time series.
In this scenario, the canonical and the microcanonical ensembles are equivalent if
$\lim_{n\to\infty}\frac{D_{\text{KL}}(\mathbf{c}^{*})}{n}=\lim_{n\to\infty}\frac{\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\rvert}{n}=0$ (35)
@@
$D_{\text{KL}}(\mathbf{c}^{*})=\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\rvert=o(n).$ (36)
@@
as it follows directly from (III.3). Thus, if the two ensembles are equivalent, the difference in description lengths between the corresponding models will be $o(n)$. However, a similar conclusion cannot be immediately derived when non-equivalence holds and $D_{\text{KL}}$ grows at least linearly with $n$. Indeed, the likelihood and complexity terms in (III.3) are of the same order and have opposite signs; therefore, evaluating the order of $\Delta\text{DL}$ is a non trivial task in this case. The following section addresses this problem by explicitly evaluating $\Delta\text{DL}$ in two notable examples of maximum-entropy models.

### v3

@@ Abstract
The (non-)equivalence of canonical and microcanonical ensembles is a central concept in statistical physics. Non-equivalence has recently been established in models with an extensive number of constraints, common, e.g., in network science, via a non-vanishing difference in relative entropy, corresponding to higher microcanonical log-likelihood per node. However, from a model selection perspective, comparing canonical and microcanonical models requires consideration of both log-likelihood and complexity. To compare both terms under the Minimum Description Length (MDL) principle, we compute the Normalized Maximum Likelihood (NML) of binary canonical and microcanonical models, finding that (i) microcanonical models, though higher in likelihood, are always more complex, making the choice of model non-trivial. (ii) The optimal model choice depends on the empirical values of the constraints: the canonical model performs best when its fit to observed data exceeds its uniform average fit across all data. (iii) Notably, in the thermodynamic limit the difference in description length per node vanishes for the equivalent models considered but persists otherwise, showing that non-equivalence implies extensive differences between large canonical and microcanonical models. Finally, we compare the NML approach to Bayesian methods, showing that (iv) the choice of priors, while practically uninfluential in equivalent models, becomes crucial when an extensive number of constraints is enforced, possibly leading to very different outcomes.
@@ I Introduction
Entropy maximization provides a principled approach to inference, resulting in maximally random models under specified constraints. These models have found applications across a wide range of scientific disciplines, such as network science [1, 2, 3] and time-series analysis [4, 5].
@@
Two distinct formulations of maximum-entropy models are possible depending on how constraints are enforced. Canonical models are defined by requiring the value of the constraints to be satisfied on average. Microcanonical models are defined by requiring the value of the constraints to be satisfied exactly. These two formulations do not necessarily lead to the same description of a given system, i.e., they are, in general, non-equivalent.
@@
Non-equivalence between canonical and microcanonical models can be formally defined in several (under mild conditions, equivalent) ways [6]. A particularly convenient definition is the measure-level one, which is rooted in information theory and involves the Kullback-Leibler (KL) divergence between the microcanonical and canonical probability distributions. According to this definition, (non-)equivalence is characterized by a (non-)vanishing relative entropy density as the system size approaches infinity. Interestingly, since the KL divergence coincides with the difference between the canonical and the microcanonical log-likelihood functions, measure-level (non-)equivalence can be inspected by simply considering the asymptotic behavior of such difference [7]. Non-equivalence has been recently shown to arise in models that involve an extensive number of constraints (e.g.,the Configuration Model for networks [7, 8, 9]) and may have practical implications for the expected macroscopic properties of the system under study (see [10] for an example concerning bipartite networks).
@@
Whenever the two approaches are equivalent, they identify the same set of typical configurations. In these cases, preferring one formulation over the other becomes a matter of convenience (e.g.,computational efficiency). Whenever the two approaches are non-equivalent, the choice depends on the specific characteristics of the system under study and the nature of the constraints involved. For example, in the context of choosing a null model to be compared with real data for, e.g., pattern detection, it has been argued that, when non-equivalence holds, the ensemble choice should be based on a theoretical guiding principle [11, 9, 12, 13, 14]. In this case, if the data are expected to be noisy, meaning that the measured values of the constraints generally do not match the true (unobservable) ones, one should prefer the canonical version of the null model because it allows for fluctuations in the constraint quantities, while the microcanonical one would paradoxically assign zero probability to the true configuration of the system. Conversely, if, by hypothesis, the measured values of the constraints are not affected by error or noise, the microcanonical model should be preferred.
@@
When there is no prior expectation available about the presence or absence of noise in the empirical values of the constraints, the decision between hard and soft constraints should be based on posterior evidence, i.e., which of the two null models best describes the data. This immediately leads to a problem of model selection between canonical and microcanonical ensembles.
@@
Information theory provides a powerful framework for model selection by offering criteria that capture the trade-off between a model ‘goodness of fit,’ quantified by its log-likelihood, and a model ‘complexity.’ The two historically most popular model selection criteria are the Akaike information criterion (AIC) [15] and the Bayesian information criterion (BIC) [16], according to which one should pick the model minimizing
@@
respectively, where $\mathcal{L}$ is the model likelihood, $k$ is the number of its parameters and $V$ is the sample size. Canonical and microcanonical models defined by the same set of constraints share the same number of parameters. Therefore, according to both AIC and BIC, their complexity is the same, and a comparison based on these criteria would reduce to a mere comparison between their log-likelihood functions. This would naively lead to the conclusion that the microcanonical model is always the best-scoring one since it has the highest likelihood. However, neither AIC nor BIC can be computed for microcanonical models since they are derived under regularity assumptions that are not even remotely met by the latter. Moreover, both criteria are asymptotic results derived under the assumption of a finite number of parameters as the system size approaches infinity [17]. This assumption is not met by models showing non-equivalence, whose number of parameters scales with the size of the system.
@@
To overcome these issues, here we consider the Minimum Description Length (MDL) principle [18, 19, 20], a family of information theoretical approaches to model selection seeking the model that provides the shortest description of data. Specifically, we focus on the approach based upon the Normalized Maximum Likelihood (NML), or Shtarkov, distribution [21], our aim being that of investigating the interplay between the difference of log-likelihoods and complexities in light of ensemble non-equivalence. Indeed, while the impact of non-equivalence on the difference between log-likelihoods is quite well understood, its effects on the difference between complexities remain an unexplored topic within this framework. To encompass a broad spectrum of applications, our study focuses on maximum-entropy models suitable for studying systems represented by binary rectangular matrices (e.g., bipartite networks, time series). More precisely, we aim at determining i) whether the influence of non-equivalence persists even when the complexity term is accounted for and ii) if non-equivalent canonical and microcanonical models should, in fact, be considered as different from a model selection perspective. We conclude that the answer to both questions is yes, raising the question of whether non-equivalence can be defined on a model selection level. Moreover, as a consequence of our results, we show that, in a Bayesian context, different priors can result in vastly different description lengths when an extensive number of constraints is enforced.
@@
The rest of the paper is organized as follows. Section II introduces MDL for discrete data and the Normalized Maximum Likelihood (NML) approach to define description lengths. In Section III, we provide a general expression for the difference between the description length of canonical and microcanonical models as a function of the Kullback-Leibler divergence between the two distributions. In Section IV, we explicitly derive and compare the description length of both the canonical and the microcanonical formulations of models defined by one global constraint (i.e., the sum of the entries of a matrix) as well as one-sided local constraints (i.e., the row-specific sums of a matrix). While the first constraint leads to ensemble equivalence, the second set of constraints leads to ensemble non-equivalence. Finally, Section V reviews the relationship between the NML-based and Bayesian approaches when non-equivalence holds.
@@ II The Minimum Description Length principle
However, the distribution $\mathcal{L}_{\mathcal{M}}(\bm{x})$ is affected by (at least) two problems having the same origin: the parameters appearing in its definition are evaluated for each individual data in $\mathcal{X}$. As a consequence, i) $\mathcal{L}_{\mathcal{M}}(\bm{x})$ cannot be a probability distribution with support in $\mathcal{X}$ as it is not properly normalized, and ii) adopting $\mathcal{L}_{\mathcal{M}}(\bm{x})$ may lead to severe overfitting since it is defined a posteriori relative to the observations.
To solve these problems, one can build a so-called universal distribution relative to $\mathcal{M}$, i.e.,a unique, representative distribution $\bar{P}_{\mathcal{M}}(\bm{x})$ of the statistical model $\mathcal{M}$. Once equipped with a universal distribution, the description length of $\bm{x}$ through is $\mathcal{M}$ is defined as
@@
Universal distributions (which don’t necessarily belong to $\mathcal{M}$) can be defined in several ways. In what follows, we will consider the MDL recipe based upon the NML universal distribution (for a more formal introduction to universal coding, we redirect the interested reader to [18]).
@@ II.1 The Normalized Maximum Likelihood universal distribution
The NML distribution corresponds to the distribution attaining the minimum point-wise regret in the worst-case scenario [18], i.e.
@@
In words, it is the distribution that a priori requires the minimum number of nats when compared to the shortest description length a posteriori, that is $-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))$. The minimum is found in the worst-case scenario, i.e.,for the data $\bm{x}$ for which this difference is maximal. The unique solution to this minimax problem reads
@@
is named parametric complexity of $\mathcal{M}$. Notice that $\mathcal{X}$ represents the set of all possible data sets of fixed size $V$, i.e.,all the observable samples of size $V$. Thus, $\text{COMP}_{\mathcal{M}}$ depends on both the model and $\mathcal{X}$. Finally, we can rewrite equation (II.1) as
@@ III Description length of canonical and microcanonical models
Although computing the complexity term within the NML-based description length can be challenging, models defined by a sufficient statistic admit a simpler expression for this term. This section introduces an important class of models of this kind, the maximum-entropy models (MEMs), both in their canonical and microcanonical formulations.
Specifically, we will focus on discrete data that can be represented as a binary $n\times m$ matrix $\mathbf{G}$, i.e.
@@
and denote by $\mathcal{G}$ the set of all such matrices that, when endowed with a probability distribution $P(\mathbf{G})$, constitutes a proper ensemble of matrices. The matrix representation is often used to describe systems composed of $n$ elements to be modelled (corresponding to the number of rows), each characterized by $m$ state variables or degrees of freedom (corresponding to the number of columns). The size of the data set is fixed and supposed to represent the number of independent entries of the matrix, i.e.,$nm$.
@@
while constraining the sufficient statistics. If the model is required to reproduce the value of the constraints exactly, i.e.,$\mathbf{c}(\mathbf{G})=\mathbf{c}^{*}$, one obtains a microcanonical model. If, instead, the model is required to reproduce the value of the constraints on average, i.e.,$\langle\mathbf{c}(\mathbf{G})\rangle_{P}=\mathbf{c}^{*}$ with $\langle\cdot\rangle_{P}$ being the ensemble average induced by distribution $P$, one obtains the corresponding canonical formulation. This approach, introduced by Gibbs [22] and reprised by Jaynes [23], leads to ensembles of matrices that reproduce (either exactly or on average) the constrained quantities while randomizing everything else.
@@
with $\delta_{i,j}$ being the Kronecker delta. As this set may be empty for some values of $\mathbf{c}$, we will solely focus on the graphical values, i.e.,the set $\mathcal{C}=\{\mathbf{c}:\Omega(\mathbf{c})\neq\emptyset\}$ of values that are verified by at least one configuration, and denote it with $\mathcal{N}_{\mathcal{C}}=|\mathcal{C}|$ its cardinality. In other words, the sum in equation (14) runs over the graphical values of the sufficient statistics rather than over all possible data.
@@ III.1 Description length of microcanonical models
$P_{\textrm{mic}}(\mathbf{G};\mathbf{c})=\begin{cases}\frac{1}{\Omega(\mathbf{c})}&\text{if}\>\mathbf{c}(\mathbf{G})=\mathbf{c}\\ 0&\text{else},\end{cases}$ (16)
@@
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{c})=\ln\frac{1}{\Omega(\mathbf{c})}=-\ln\Omega(\mathbf{c}).$ (17)
@@ III.2 Description length of canonical models
In fact, it can be shown [8, 7] that
@@ III.3 Comparing canonical and microcanonical models
i.e.,the canonical likelihood is always smaller than or equal to the microcanonical one. For the same reason, it is straightforward to see that the canonical complexity is always smaller than, or equal to, the microcanonical one:
@@
where $Q_{\text{can}}^{*}\equiv Q_{\text{can}}(\mathbf{c}^{*},\bm{\theta}^{*})$. This expression depends only on the canonical model, providing a new interpretation of the comparison between the canonical and the microcanonical description length: the canonical model provides a shorter description length whenever $\Delta\text{DL}(\mathbf{G}^{*})<0$, i.e.,whenever $Q_{\text{can}}^{*}>\bar{Q}_{\text{can}}$. In other words, the canonical model is the best-scoring model whenever the maximum likelihood it assigns to the observed value $\mathbf{c}^{*}$ is higher than the maximum likelihood averaged over all possible values of the sufficient statistic (or, more intuitively, if it performs ‘better than on average’ on the observed sufficient statistic). Remarkably, this is in open contrast with a comparison based only on the likelihoods, according to which the microcanonical formulation should always be preferred, independently of the observed data.
Expression (III.3) shows that the entity of the difference between the two description lengths crucially depends on the value of the KL divergence. As said in the introduction, the canonical and the microcanonical ensembles are (non-)equivalent if the relative $\text{D}_{\text{KL}}$ is (non-)vanishing as the system size approaches infinity. When studying matrices, identifying the system size with the number of entries, i.e.,$nm$, may be inappropriate as adding (only) one entry to a matrix without altering its structure is, in fact, impossible. In these cases, it may be much more reasonable to identify the system size with the number $n$ of items: increasing the system size from $n$ to $n+1$ would, thus, correspond to adding a node to a network or a time-point to a time series.
In this scenario, the canonical and the microcanonical ensembles are equivalent in the measure sense [6] if
$\lim_{n\to\infty}\frac{D_{\text{KL}}(\mathbf{c}^{*})}{n}=\lim_{n\to\infty}\frac{\left\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\right\rvert}{n}=0$ (35)
@@
$D_{\text{KL}}(\mathbf{c}^{*})=\left\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\right\rvert=o(n).$ (36)
@@
as it follows directly from (III.3). Thus, if the two ensembles are equivalent, the difference in description lengths between the corresponding models will be $o(n)$. However, a similar conclusion cannot be immediately derived when non-equivalence holds and $D_{\text{KL}}$ grows at least linearly with $n$. Indeed, the likelihood and complexity terms in (III.3) are of the same order and have opposite signs; therefore, evaluating the order of $\Delta\text{DL}$ is a non-trivial task in this case. The following section addresses this problem by explicitly evaluating $\Delta\text{DL}$ in two notable examples of maximum-entropy models.